# CtxMIM: Context-Enhanced Masked Image Modeling for Remote Sensing Image Understanding

### v3

@@ Abstract
keywords:
Self-supervised learning , masked image modeling , Siamese , context consistency constraint , remote sensing
journal: ISPRS Journal of Photogrammetry and Remote Sensing
[…]
1 Introduction
Remote sensing images have been widely used in many applications, including precision agriculture, disaster management, city planning, and environment monitoring. With a growing number of satellites collecting remote sensing images ceaselessly, abundant remote sensing images are easy to access. However, annotations are technically challenging due to highly time-consuming and expertise knowledge than general benchmark datasets, e.g., ImageNet (Russakovsky et al., 2015). Recently, self-supervised learning (SSL) has been increasingly emerging to pretrain models on extensive unlabeled data and demonstrated more robust performance than supervised learning on classification, segmentation, and detection. Therefore, the SSL paradigm has been receiving much attention from the remote sensing community and exploring to obtain an efficient pretrained model on large-scale unlabeled data for remote sensing image understanding.
@@ 1 Introduction
Figure 2: Reconstructive results on remote sensing images. The first column is raw images. The other triples show the masked image with different masking ratios (left), SimMIM (Xie et al., 2022) reconstruction (middle), and our CtxMIM reconstruction (right). SimMIM (Xie et al., 2022) and the proposed CtxMIM are pretrained on our collected dataset. Although more and more land objects are masked as the ratio becomes larger, our CtxMIM can still reconstruct masked images better in terms of texture and content.
Existing advanced SSL methods (Manas et al., 2021; Ayush et al., 2021; Mall et al., 2023; Huang et al., 2024) primarily adopt contrastive learning (Chen et al., 2020b, a; Chen and He, 2021) for self-supervised remote sensing representation learning. Indeed, contrastive approaches learn representations invariant to season, geographical location, illumination, and viewing angle by employing extensive data augmentation, which have exceeded supervised learning-based methods. However, state-of-the-art contrastive methods typically require careful pretraining data organization in terms of temporal and geographical dimensions to prevent the pretraining collapse. Additionally, these methods generally concentrate on image-level feature learning by comparing two augmented views.
Recently, SSL based on masked image modeling (MIM) has demonstrated potential performance in natural image representation learning (He et al., 2022a; Xie et al., 2022). These MIM methods reconstruct masked patches by randomly masking an image with a high masking ratio (e.g., 75% in MAE (He et al., 2022a) and 80% in SimMIM (Xie et al., 2022)), which can learn good representations without extensive data augmentation. The pixel reconstruction idea of MIM is well-suited for remote sensing tasks since remote sensing analysis typically requires object-level or pixel-level features with fine-grained details rather than image-level features. However, applying MIM from natural images to remote sensing images still has limitations due to the considerably high object density in remote sensing images. Natural images have apparent foreground objects in a relatively simple background, e.g., the cat, castle, and ambulance at the top of Figure 1, which can be easily repaired by inferring the whole from some parts through rich contextual information. Since remote sensing images are captured by sensors mounted on satellites, they generally contain multiple objects in a vast and complicated scene and typically have no apparent foreground objects, as illustrated at the bottom of Figure 1. Unlike natural images, some small land objects may be entirely masked by randomly masking, making pretraining convergence difficult, as demonstrated in ablation studies of subsection 4.6. Hence, MIM methods with a high masking ratio would hamper semantic extrapolating for remote sensing representation learning due to missing contextual information, as shown in Figure 2.
@@
We evaluate our approach on different tasks, including land cover classification, semantic segmentation, object detection, and instance segmentation. Following SOTA remote sensing SSL methods, we use EuroSAT (Helber et al., 2019), NWPU-RESISC45 (Cheng et al., 2017), DOTA (Xia et al., 2018), and SpaceNet (Van Etten et al., 2018) for comparison. Extensive experiments demonstrate that our method outperforms SOTA SSL methods and is more effective for remote sensing tasks than supervised learning on the common ImageNet (Russakovsky et al., 2015). In summary, our contributions are:
@@
The remainder of this paper is organized as follows: Section 2 reviews related studies. Section 3 introduces our proposed CtxMIM in detail. Section 4 first describes experimental settings, comparison methods, and evaluation metrics, then reports and discusses the experimental results quantitatively and qualitatively. Section 5 analyzes remote sensing pretraining in terms of model structure, performance, limitations, and future work. Section 6 concludes this work.
2 Related work
@@
In this section, we first introduce two dominant self-supervised learning frameworks (i.e., contrastive learning, and reconstructive learning) in the computer vision field. Motivated by increasing success in many computer vision tasks, self-supervised learning has been a prevalent paradigm for remote sensing pretraining, which is discussed in the following 2.3.
2.1 Contrastive Learning
@@
Contrastive learning based methods (Oord et al., 2018; Tian et al., 2020) learn good representations by minimizing the distance between representations of positive pairs augmented from one image and pushing representations of negative pairs from different images away. SimCLR (Chen et al., 2020a) systematically analyzes the principles of contrastive learning methods and designs a simple framework for representation learning based on a Siamese network and extensive data augmentations. MoCos (He et al., 2020; Chen et al., 2020b) introduce a dynamic large queue from different batches for contrastive learning and employ a moving-averaged encoder to maintain consistency. Since these methods require large batches or negative pairs to avoid collapsing, some methods (Grill et al., 2020; Caron et al., 2020; Chen and He, 2021; Zbontar et al., 2021; Li et al., 2021c) redesign the network structures, propose an innovative objective function, or introduce online clustering for representation learning. Recent contrastive methods (Kim and Song, 2021; Guo et al., 2022; Zhang et al., 2022; Reiss and Hoshen, 2023) have demonstrated superiority in different downstream tasks, even outperforming supervised learning. Although increasingly successful in many computer vision tasks, contrastive methods rely on extensive data augmentations and focus on image-level features, limiting the transferring performance in object-level or pixel-level cases.
2.2 Reconstructive Learning
@@
Based on the idea of masked language modeling in NLP, reconstructive learning (Bao et al., 2021; He et al., 2022a; Xie et al., 2022; Wei et al., 2022; Liu et al., 2023b) learns representations through solving the reconstructive pre-text task. This line of work randomly masks a portion of image patches using a special [MASK] token and then predicts raw pixel values using masked image modeling. SimMIM (Xie et al., 2022) only uses a lightweight prediction head to reconstruct raw pixel values of masked patches, facilitating good representation learning of the encoder. MAE (He et al., 2022a) first handles visible patches in the encoder, and then encoded patches and mask tokens are processed in the decoder, which can significantly reduce the computation overhead of the encoder. MixMAE (Liu et al., 2023a) reconstructs two original images from the mixed image by replacing the original masked patches with visible patches of another image, improving pretraining efficiency and consistency. With SOTA performance on various tasks, many related works have been studied, such as VideoMAE (Tong et al., 2022), MultiMAE (Bachmann et al., 2022), CIM (Fang et al., 2022), MSN (Assran et al., 2022), and SiamMAE (Gupta et al., 2024). However, these works have yet to be explored on large-scale remote sensing datasets.
2.3 Self-Supervised Learning in Remote Sensing
@@
Recently, self-supervised learning has been broadly applied in remote sensing domains. For instance, many methods (Kang et al., 2020; Vincenzi et al., 2021; Li et al., 2021a; Stojnic and Risojevic, 2021; Jung et al., 2021) utilize contrastive learning to learn representations. SeCo (Manas et al., 2021) contrastively learns seasonal invariance by constructing a remote sensing dataset of different temporal information. Some methods (Ayush et al., 2021; Li et al., 2021b, 2022b) consider that geographical knowledge is the unique characteristic of remote sensing images, which can be used for contrastive learning. GeRSP (Huang et al., 2024) simultaneously learns representations with general and special knowledge from remote sensing and natural images by a teacher-student architecture. It designs self-supervised and supervised pretraining branches to generate the powerful pretrained model for remote sensing image understanding. Another attempt (Cong et al., 2022; He et al., 2023b; Mendieta et al., 2023; Hong et al., 2024) explores the masked image modeling paradigm for remote sensing feature pretraining by repairing masked images. SatMAE++ (Noman et al., 2024) leverages the scale information of remote sensing imagery to pretrain a transformer framework, finally achieving representations with more scales. ScaleMAE (Reed et al., 2023) reconsiders the scale information by explicitly learning relationships between data at different scales. After encoding a masked image, Scale-MAE reconstructs low/high frequency images at lower/higher scales. CrossScaleMAE (Tang et al., 2024), based on the Masked Auto-Encoder (MAE), adopts scale augmentation techniques and introduces cross-scale consistency constraints to pretrain consistent and meaningful representations for downstream tasks. Besides, more relevant works have been studied for particular tasks, including change detection (Chen et al., 2023; de Gélis et al., 2023; Fang et al., 2023), land cover classification (Scheibenreif et al., 2022; Zeng and Geng, 2022; Bai et al., 2023; Xu et al., 2024), semantic segmentation (Li et al., 2022a; Muhtar et al., 2022; He et al., 2022b; Tang et al., 2023), and image fusion (He et al., 2023a; Sun et al., 2023; Wang et al., 2023).
@@ 3.1 Overview
SSL, a prevalent paradigm in natural images, has been applied to large unlabeled remote sensing images and learned useful representations for downstream tasks (classification, segmentation, and detection). However, SSL in remote sensing still has been limited and lagged behind that in natural images since the object density is different between natural and remote sensing images. Natural images, such as the general benchmark dataset ImageNet (Russakovsky et al., 2015), are captured in daily scenes, and objects are mostly large in size. Hence, natural images typically have apparent foreground objects in a relatively simple background, like the fox in Figure 3. Remote sensing images are captured by sensors mounted on satellites and contain multiple objects in a vast and complicated scene, especially no apparent foreground objects, as illustrated in the bottom of Figure 3. Therefore, the high object density poses challenges to the effectiveness of the SSL paradigm in remote sensing understanding, manifesting in the following issues: (1) Mismatched positive pairs when augmenting positive pairs from one image for contrastive learning, as shown in column (a) of Figure 3. (2) Missing contextual information caused by random masking strategies in reconstructive learning, as exemplified in column (b) of Figure 3.
@@
To address this issue, we propose CtxMIM, a novel context-enhanced self-supervised representation learning method based on reconstructive learning. As shown in Figure 4, CtxMIM designs a reconstructive branch and a context-enhanced generative branch using an asymmetric Siamese network, which can learn representations by masked image modeling. For the reconstructive branch, CtxMIM first uniformly splits an input $x$ into non-overlapping image patches $x^{p}$ and then randomly mask $x^{p}$ on patch-level to obtain a new input patches $\hat{x}^{p}$. Subsequently, CtxMIM takes as input $\hat{x}^{p}$ to reconstruct the raw pixel values of the masked patches through the encoder-decoder architecture. The reconstructive branch is detailed in the subsection 3.2. Unlike natural images, the object density of remote sensing images is highly uneven, resulting in missing contextual information when randomly masking image patches. To tackle this issue, we represent the original image patches as reconstructive templates rather than directly removing the masked patches. Based on this analysis, we introduce a context-enhanced generative branch to facilitate semantic reasoning by providing meaningful contextual information during the reconstruction.
@@
Moreover, the context-enhanced generative branch shares the encoder and the decoder with the reconstructive branch, which is detailed in the subsection 3.3. For pretraining, we minimize the distance between raw pixel values x and the reconstructive values y for reconstructive and context-enhanced branches. Additionally, a context consistency loss is proposed to guide the reconstructive process of masked patches, which is described in subsection 3.4. Finally, our proposed CtxMIM is pretrained on a large-scale unlabeled dataset and learns impressive representations, which is beneficial for various downstream tasks (e.g., image-level, object-level, and pixel-level).
3.2 Reconstruction of masked image patches
@@
In the reconstruction stage, CtxMIM aims to learn local fine-grained features by restoring masked image patches. Given an input image $x\in\mathbb{R}^{C\times H\times W}$, CtxMIM firstly reshapes $x$ to get image sequence patches $x^{p}\in\mathbb{R}^{N\times P^{2}C}$, where $H,W,C$ are the image height, width, and channel, respectively, $P$ is the size of each image patch (i.e., the height and width), and $N=(H/P)\cdot(W/P)$ is the number of image patches. Then, a patch-wise masking operation randomly masks some image patches, and a patch embedding operation encodes each image patch to get the new sequence input $\hat{x}^{p}$. Subsequently, an encoder $f_{\theta}$ takes the sequence input $\hat{x}^{p}$ to extract latent representations $\hat{h}^{p}$, which are used in downstream tasks after pretraining. Finally, a decoder $g_{\theta}$ is used to reconstruct pixel values of masked patches $\hat{y}^{p}$ on the latent representations. In our paper, we take Swin Transformer (Liu et al., 2021) as the feature encoder $f_{swin}$ and a lightweight prediction head as the decoder $g_{mlp}$ to output raw pixel values of the masked patches by following SimMIM (Xie et al., 2022), which can be formulated as:
@@
3.3 Context-enhanced generative branch
@@ 3.4 Training objective
The proposed CtxMIM adopts Swin-B as the encoder and is implemented by the PyTorch framework. We use an AdamW optimizer with hyper-parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-8}$, and the learning rate is set to $1\times 10^{-5}$. The training lasts 200 epochs with 8 NVIDIA Tesla V100 GPUs. For pretraining, we collect a large-scale unlabeled remote sensing dataset captured from WorldView-3 using Google Earth Engine (Gorelick et al., 2017). This pretraining dataset is characterized by diverse topography and terrain, covering most regions (cities, villages, rivers, mountains, woodlands, and other terrain) of Asia.
4 Experiments

### v4

@@ Abstract
Index Terms:
Self-supervised learning, masked image modeling, Siamese, context consistency constraint, remote sensing.
I Introduction
Remote sensing images have been widely used in many applications, including precision agriculture, disaster management, city planning, and environment monitoring. With a growing number of satellites collecting remote sensing images ceaselessly, abundant remote sensing images are easy to access. However, annotations are technically challenging due to highly time-consuming and expertise knowledge than general benchmark datasets, e.g., ImageNet [1]. Recently, self-supervised learning (SSL) has been increasingly emerging to pretrain models on extensive unlabeled data and demonstrated more robust performance than supervised learning on classification, segmentation, and detection. Therefore, the SSL paradigm has been receiving much attention from the remote sensing community and exploring to obtain an efficient pretrained model on large-scale unlabeled data for remote sensing image understanding.
@@ I Introduction
Figure 2: Reconstructive results on remote sensing images. The first column is raw images. The other triples show the masked image with different masking ratios (left), SimMIM [2] reconstruction (middle), and our CtxMIM reconstruction (right). SimMIM [2] and the proposed CtxMIM are pretrained on our collected dataset. Although more and more land objects are masked as the ratio becomes larger, our CtxMIM can still reconstruct masked images better in terms of texture and content.
Existing advanced SSL methods [3, 4, 5] primarily adopt contrastive learning [6, 7, 8] for self-supervised remote sensing representation learning. Indeed, contrastive approaches learn representations invariant to season, geographical location, illumination, and viewing angle by employing extensive data augmentation, which have exceeded supervised learning-based methods. However, state-of-the-art contrastive methods typically require careful pretraining data organization in terms of temporal and geographical dimensions to prevent the pretraining collapse. Additionally, these methods generally concentrate on image-level feature learning by comparing two augmented views.
Recently, SSL based on masked image modeling (MIM) has demonstrated potential performance in natural image representation learning [9, 2]. These MIM methods reconstruct masked patches by randomly masking an image with a high masking ratio (e.g., 75% in MAE [9] and 80% in SimMIM [2]), which can learn good representations without extensive data augmentation. The pixel reconstruction idea of MIM is well-suited for remote sensing tasks since remote sensing analysis typically requires object-level or pixel-level features with fine-grained details rather than image-level features. However, applying MIM from natural images to remote sensing images still has limitations due to the considerably high object density in remote sensing images. Natural images have apparent foreground objects in a relatively simple background, e.g., the cat, castle, and ambulance at the top of Figure 1, which can be easily repaired by inferring the whole from some parts through rich contextual information. Since remote sensing images are captured by sensors mounted on satellites, they generally contain multiple objects in a vast and complicated scene and typically have no apparent foreground objects, as illustrated at the bottom of Figure 1. Unlike natural images, some small land objects may be entirely masked by randomly masking, making pretraining convergence difficult, as demonstrated in ablation studies of subsection IV-F. Hence, MIM methods with a high masking ratio would hamper semantic extrapolating for remote sensing representation learning due to missing contextual information, as shown in Figure 2.
@@
We evaluate our approach on different tasks, including land cover classification, semantic segmentation, object detection, and instance segmentation. Following SOTA remote sensing SSL methods, we use EuroSAT [10], NWPU-RESISC45 [11], DOTA [12], and SpaceNet [13] for comparison. Extensive experiments demonstrate that our method outperforms SOTA SSL methods and is more effective for remote sensing tasks than supervised learning on the common ImageNet [1]. In summary, our contributions are:
@@
The remainder of this paper is organized as follows: Section II reviews related studies. Section III introduces our proposed CtxMIM in detail. Section IV first describes experimental settings, comparison methods, and evaluation metrics, then reports and discusses the experimental results quantitatively and qualitatively. Section V analyzes remote sensing pretraining in terms of model structure, performance, limitations, and future work. Section VI concludes this work.
II Related work
@@
In this section, we first introduce two dominant self-supervised learning frameworks (i.e., contrastive learning, and reconstructive learning) in the computer vision field. Motivated by increasing success in many computer vision tasks, self-supervised learning has been a prevalent paradigm for remote sensing pretraining, which is discussed in the following II-C.
II-A Contrastive Learning
@@
Contrastive learning based methods [14, 15] learn good representations by minimizing the distance between representations of positive pairs augmented from one image and pushing representations of negative pairs from different images away. SimCLR [7] systematically analyzes the principles of contrastive learning methods and designs a simple framework for representation learning based on a Siamese network and extensive data augmentations. MoCos [16, 6] introduce a dynamic large queue from different batches for contrastive learning and employ a moving-averaged encoder to maintain consistency. Since these methods require large batches or negative pairs to avoid collapsing, some methods [17, 18, 8, 19] redesign the network structures, propose an innovative objective function, or introduce online clustering for representation learning. Recent contrastive methods [20, 21, 22, 23] have demonstrated superiority in different downstream tasks, even outperforming supervised learning. Although increasingly successful in many computer vision tasks, contrastive methods rely on extensive data augmentations and focus on image-level features, limiting the transferring performance in object-level or pixel-level cases.
II-B Reconstructive Learning
@@
Based on the idea of masked language modeling in NLP, reconstructive learning [9, 2, 24] learns representations through solving the reconstructive pre-text task. This line of work randomly masks a portion of image patches using a special [MASK] token and then predicts raw pixel values using masked image modeling. SimMIM [2] only uses a lightweight prediction head to reconstruct raw pixel values of masked patches, facilitating good representation learning of the encoder. MAE [9] first handles visible patches in the encoder, and then encoded patches and mask tokens are processed in the decoder, which can significantly reduce the computation overhead of the encoder. MixMAE [25] reconstructs two original images from the mixed image by replacing the original masked patches with visible patches of another image, improving pretraining efficiency and consistency. With SOTA performance on various tasks, many related works have been studied, such as VideoMAE [26], MultiMAE [27], DMDSC [28], MCGMAE [29], and SiamMAE [30].
II-C Self-Supervised Learning in Remote Sensing
@@
Recently, self-supervised learning has been broadly applied in remote sensing domains. For instance, many methods [31, 32, 33, 34, 35] utilize contrastive learning to learn representations. SeCo [3] contrastively learns seasonal invariance by constructing a remote sensing dataset of different temporal information. Some methods [4, 36, 37] consider that geographical knowledge is the unique characteristic of remote sensing images, which can be used for contrastive learning. GeRSP [38] simultaneously learns representations with general and special knowledge from remote sensing and natural images by a teacher-student architecture. It designs self-supervised and supervised pretraining branches to generate the powerful pretrained model for remote sensing image understanding. Another attempt [39, 40, 41, 42] explores the masked image modeling paradigm for remote sensing feature pretraining by repairing masked images. SatMAE++ [43] leverages the scale information of remote sensing imagery to pretrain a transformer framework, finally achieving representations with more scales. ScaleMAE [44] reconsiders the scale information by explicitly learning relationships between data at different scales. After encoding a masked image, Scale-MAE reconstructs low/high frequency images at lower/higher scales. CrossScaleMAE [45], based on the Masked Auto-Encoder (MAE), adopts scale augmentation techniques and introduces cross-scale consistency constraints to pretrain consistent and meaningful representations for downstream tasks. Besides, more relevant works have been studied for particular tasks, including change detection [46, 47, 48], land cover classification [49, 50, 51], and semantic segmentation [52, 53, 54].
@@ III-A Overview
SSL, a prevalent paradigm in natural images, has been applied to large unlabeled remote sensing images and learned useful representations for downstream tasks (classification, segmentation, and detection). However, SSL in remote sensing still has been limited and lagged behind that in natural images since the object density is different between natural and remote sensing images. Natural images, such as the general benchmark dataset ImageNet [1], are captured in daily scenes, and objects are mostly large in size. Hence, natural images typically have apparent foreground objects in a relatively simple background, like the fox in Figure 3. Remote sensing images are captured by sensors mounted on satellites and contain multiple objects in a vast and complicated scene, especially no apparent foreground objects, as illustrated in the bottom of Figure 3. Therefore, the high object density poses challenges to the effectiveness of the SSL paradigm in remote sensing understanding, manifesting in the following issues: (1) Mismatched positive pairs when augmenting positive pairs from one image for contrastive learning, as shown in column (a) of Figure 3. (2) Missing contextual information caused by random masking strategies in reconstructive learning, as exemplified in column (b) of Figure 3.
@@
To address this issue, we propose CtxMIM, a novel context-enhanced self-supervised representation learning method based on reconstructive learning. As shown in Figure 4, CtxMIM designs a reconstructive branch and a context-enhanced generative branch using an asymmetric Siamese network, which can learn representations by masked image modeling. For the reconstructive branch, CtxMIM first uniformly splits an input $x$ into non-overlapping image patches $x^{p}$ and then randomly mask $x^{p}$ on patch-level to obtain a new input patches $\hat{x}^{p}$. Subsequently, CtxMIM takes as input $\hat{x}^{p}$ to reconstruct the raw pixel values of the masked patches through the encoder-decoder architecture. The reconstructive branch is detailed in the subsection III-B. Unlike natural images, the object density of remote sensing images is highly uneven, resulting in missing contextual information when randomly masking image patches. To tackle this issue, we represent the original image patches as reconstructive templates rather than directly removing the masked patches. Based on this analysis, we introduce a context-enhanced generative branch to facilitate semantic reasoning by providing meaningful contextual information during the reconstruction.
@@
Moreover, the context-enhanced generative branch shares the encoder and the decoder with the reconstructive branch, which is detailed in the subsection III-C. For pretraining, we minimize the distance between raw pixel values x and the reconstructive values y for reconstructive and context-enhanced branches. Additionally, a context consistency loss is proposed to guide the reconstructive process of masked patches, which is described in subsection III-D. Finally, our proposed CtxMIM is pretrained on a large-scale unlabeled dataset and learns impressive representations, which is beneficial for various downstream tasks (e.g., image-level, object-level, and pixel-level).
III-B Reconstruction of masked image patches
@@
In the reconstruction stage, CtxMIM aims to learn local fine-grained features by restoring masked image patches. Given an input image $x\in\mathbb{R}^{C\times H\times W}$, CtxMIM firstly reshapes $x$ to get image sequence patches $x^{p}\in\mathbb{R}^{N\times P^{2}C}$, where $H,W,C$ are the image height, width, and channel, respectively, $P$ is the size of each image patch (i.e., the height and width), and $N=(H/P)\cdot(W/P)$ is the number of image patches. Then, a patch-wise masking operation randomly masks some image patches, and a patch embedding operation encodes each image patch to get the new sequence input $\hat{x}^{p}$. Subsequently, an encoder $f_{\theta}$ takes the sequence input $\hat{x}^{p}$ to extract latent representations $\hat{h}^{p}$, which are used in downstream tasks after pretraining. Finally, a decoder $g_{\theta}$ is used to reconstruct pixel values of masked patches $\hat{y}^{p}$ on the latent representations. In our paper, we take Swin Transformer [55] as the feature encoder $f_{swin}$ and a lightweight prediction head as the decoder $g_{mlp}$ to output raw pixel values of the masked patches by following SimMIM [2], which can be formulated as:
@@
III-C Context-enhanced generative branch
@@ III-D Training objective
The proposed CtxMIM adopts Swin-B as the encoder and is implemented by the PyTorch framework. We use an AdamW optimizer with hyper-parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-8}$, and the learning rate is set to $1\times 10^{-5}$. The training lasts 200 epochs with 8 NVIDIA Tesla V100 GPUs. For pretraining, we collect a large-scale unlabeled remote sensing dataset captured from WorldView-3 using Google Earth Engine [56]. This pretraining dataset is characterized by diverse topography and terrain, covering most regions (cities, villages, rivers, mountains, woodlands, and other terrain) of Asia.
IV Experiments